H3PO4 (aq) ⇌ H + (aq) + H2PO4 - (aq)      Ka = 7.2 x 10 -3


Give the forward reaction.


1
Expert's answer
2021-03-11T08:45:20-0500

H3PO4 (aq) ⇌ H + (aq) + H2PO4 - (aq)      Ka = 7.2 x 10 -3


Give the forward reaction.



Correct option is

B

2.92We have, Keq​=Kb​Kf​Now, Kf​=2.38×10−4;Kb​=8.15×10−5⇒Keq​=8.15×10−52.38×10−4​=2.92

Correct option is

B

28Kc=KbKf​. Now Kc = 10, Kb = 2.8 ∴kf=10×2.8∴Kf=28